1. A coil with 25 turns of wire is wrapped around a hollow tube with an area of 1.8 m². Each turn has the same area as the tube. A uniform magnetic field is applied at a right angle to the plane of the coil. If the field increases uniformly from 0.00 T to 0.55 T in 0.85 s, find the magnitude of the induced emf in the coil. If the resistance in the coil is 2.5 , find the magnitude of the induced current in the coil.



2. A coil consists of 200 turns of wire having a total resistance of 2.0. Each turn is a square of side 18 cm, and a uniform magnetic field directed perpendicular to the plane of the coil is turned on. If the field changes linearly from 0 to 0.50 Tin 0.80 s, what is the magnitude of the induced emf in the coil while the field is changing? What is the magnitude of the induced current in the coil while the field is changing?

1
Expert's answer
2022-05-31T16:11:30-0400

1.

"\varepsilon=\frac{nS∆B}{∆t}=29~V,"

"I=\frac \varepsilon R=11.6~A,"

2.

"\varepsilon=\frac{nS∆B}{∆t}=4~V,"

"I=\frac \varepsilon R=2~A."